Give me IMP derivations from Electric Current And IMP Q which Board Can be ask in exam This year 2016

Grade:12

1 Answers

Saurabh Kumar
8 years ago
I = Q/t, Q = neV/t = n e (Al)/t = naeVd
Concept of drift velocity of the mobile charge career.